OpenStudy (anonymous):

The function f(x) = 396(3)x represents the growth of a dragonfly population every year in a remote swamp. Pierre wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Pierre's purposes?

OpenStudy (anonymous):

A. f(x)= 33(3)^x b. F(x)=396(3^12)x/12 C. f(x)=396(3^1/12)^12x D. f(x)=4752(3)^x

OpenStudy (anonymous):

So, I got C?

OpenStudy (irishboy123):

**absent typos**: \(\large 396(3^{\frac{1}{12}})^{12x} = 396(3^{\frac{1}{12} \times 12x}) = 396(3)^{x}\) ie back to where you started. plug a number in and see. none of these look right https://gyazo.com/5b74d0c9b3753f89f9467f1bc3f2858c

OpenStudy (irishboy123):

screenshot?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

f(x) = 33(3)x f(x) = 396(312)the x over 12 power f(x) = 396(3 to the one twelfth power)12x f(x) = 4752(3)x

OpenStudy (campbell_st):

this is the 2nd time I have seen one of these questions and they don't make much sense to me... I'd expect for this exponential model if the time changes to months the value of x needs to be divided by 12 so you can test by looking at the model \[f(1) = 396(3)^{\frac{1}{12}}\] then find f(2) and f(3) to see if there is growth hope it helps

OpenStudy (anonymous):

That"s not an answer choice:( @campbell_st

OpenStudy (campbell_st):

well remember the index law for power of a power, multiply the powers \[(x^a)^b = x^{a \times b}\] so you have \[3^{\frac{x}{12}} = (3^?)^x\]

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

So would I be right when I said, C?